OpenStudy (anonymous):

please help me question below with answer choices | | v v Choose the equation that represents the graph below. Graph of a line passing through points negative 3 comma 0 and 0 comma 3 y = x − 3 y = −x + 3 y = −x − 3 y = x + 3

OpenStudy (anonymous):

@texaschic101 @e.cociuba @Dobby1 @DebbieG @agent0smith

ganeshie8 (ganeshie8):

where is the graph

OpenStudy (anonymous):

i actually explained how it looks like but do you prefer a picture?

ganeshie8 (ganeshie8):

nvm they gave two points already

ganeshie8 (ganeshie8):

no, its a very good explanation. its enough :)

OpenStudy (anonymous):

ok thnx

ganeshie8 (ganeshie8):

its passing thru (0, 3) that means, when x = 0, it has a value of 3 so, y-intercept = 3 it looks like below :- y = mx + 3

ganeshie8 (ganeshie8):

we just need to find m

ganeshie8 (ganeshie8):

u knw how to find slope m ? :)

OpenStudy (anonymous):

is it x1 - x2 / y1 - y2? or is that a different formula?

ganeshie8 (ganeshie8):

yes thats the formula !

ganeshie8 (ganeshie8):

apply it on given two points

ganeshie8 (ganeshie8):

3 comma 0 and 0 comma 3 (3, 0) and (0, 3) x1 y1 x2 y2

OpenStudy (anonymous):

3/-3 also -1?

ganeshie8 (ganeshie8):

yes ! good job !!

ganeshie8 (ganeshie8):

its passing thru (0, 3) that means, when x = 0, it has a value of 3 so, y-intercept = 3 it looks like below :- y = mx + 3 since we got m = -1 the equation is, y = -1x + 3

OpenStudy (anonymous):

ok so now that we know -1 is m what are we supposed to do? sorry but i'm not really good in math

OpenStudy (anonymous):

oh nvm that is the answer since we just solved correct?

ganeshie8 (ganeshie8):

yup :)

ganeshie8 (ganeshie8):

y = -1x + 3 is same as y = -x + 3

OpenStudy (anonymous):

thank you so much for the help! you and @dss are really good at explaining and helping others thank you so so so much

ganeshie8 (ganeshie8):

aww thanks too you're a great student :) wish you good luck !!

OpenStudy (anonymous):

thank you so much bye